OpenStudy (amenah8):

L'Hopitale's Rule: lim x-->0 of sinx^(cscx)

OpenStudy (amenah8):

lny=(ln(sinx))(cscx)

rvc (rvc):

proceed further using the product rule

OpenStudy (amenah8):

couldn't i also change it to ln(sinx) / (sinx) ??

OpenStudy (amenah8):

I did that and then took the derivative, but got another indeterminate.

OpenStudy (amenah8):

1/0

jimthompson5910 (jim_thompson5910):

what happens to sin(x) when x approaches 0?

OpenStudy (amenah8):

it is 0

jimthompson5910 (jim_thompson5910):

yes what happens to csc(x) when x approaches 0?

jimthompson5910 (jim_thompson5910):

be careful here: you'll need to approach 0 from both the left and right side

jimthompson5910 (jim_thompson5910):

it might help to look at a graph or table

OpenStudy (amenah8):

0

OpenStudy (amenah8):

wait, it's indeterminate.

jimthompson5910 (jim_thompson5910):

the limit isn't defined because the left hand limit and right hand limit are different agreed?

OpenStudy (amenah8):

right

jimthompson5910 (jim_thompson5910):

so overall, the limit doesn't exist or it's not defined

OpenStudy (amenah8):

so 0^infinity is the indeterminate?

jimthompson5910 (jim_thompson5910):

well that's the thing, saying \[\Large \lim_{x\to 0}\left[\csc(x)\right] = \infty\] is a false statement the limit does not exist

OpenStudy (amenah8):

so what is the indeterminate?

jimthompson5910 (jim_thompson5910):

it would have to be in these forms https://upload.wikimedia.org/math/f/0/1/f01dca26ea8fa3e4af3440e5c6a629b3.png

jimthompson5910 (jim_thompson5910):

if a piece of the limit does not exist, then the whole limit does not exist

OpenStudy (amenah8):

so is the answer simply DNE?

OpenStudy (amenah8):

Would it make a difference if it were x--> 0+

jimthompson5910 (jim_thompson5910):

I guess to be more technical, I'm using this property \[\LARGE \lim_{x \to a}\left\{\left[f(x)\right]^{g(x)}\right\} = {\lim_{x \to a}\left[f(x)\right]}^{\lim_{x \to a}\left[g(x)\right]}\]

jimthompson5910 (jim_thompson5910):

In this case, a = 0, f(x) = sin(x) and g(x) = csc(x) so, \[\LARGE \lim_{x \to 0}\left\{\left[\sin(x)\right]^{\csc(x)}\right\} = {\lim_{x \to 0}\left[\sin(x)\right]}^{\lim_{x \to 0}\left[\csc(x)\right]}\]

jimthompson5910 (jim_thompson5910):

the first limit, of sin(x) is 0 but the limit in the exponent DNE

jimthompson5910 (jim_thompson5910):

@Amenah8 `Would it make a difference if it were x--> 0+` sorry I didn't see that question til just now

jimthompson5910 (jim_thompson5910):

yes it makes a big difference because \[\Large \lim_{x \to 0^{+}}\left[\csc(x)\right]=\infty\]

jimthompson5910 (jim_thompson5910):

so, \[\LARGE \lim_{x \to 0^{+}}\left\{\left[\sin(x)\right]^{\csc(x)}\right\} = {\lim_{x \to 0^{+}}\left[\sin(x)\right]}^{\lim_{x \to 0^{+}}\left[\csc(x)\right]}\] \[\LARGE \lim_{x \to 0^{+}}\left\{\left[\sin(x)\right]^{\csc(x)}\right\} = 0^{\infty}\] which is one of the indeterminate forms

jimthompson5910 (jim_thompson5910):

actually no I apologize, it's not one of the forms. I mixed up the terms on accident

OpenStudy (amenah8):

so now I say: lny = ln(sinx)(cscx)

OpenStudy (amenah8):

so the answer is infinity

jimthompson5910 (jim_thompson5910):

\[\Large 0^{\infty} = 0\]

OpenStudy (amenah8):

oh. :)

jimthompson5910 (jim_thompson5910):

Think of it as \[\Large 0^x\] and let x approach infinity well you can simplify 0^x to get 0

OpenStudy (amenah8):

oh! i get it! thank you so much!!

jimthompson5910 (jim_thompson5910):

no problem